Find the slope and the y-intercept of the line.
10x=3y+6
My answer was the slope was 3 and the y- intercept was (0,-6) but I am not sure if that is correct? Thank you!

Answer by ewatrrr(24785)


Hi
the standard slope-intercept form for an equation of a line is y = mx + b
where m is the slope and b the y-intercept.
In order to 'read' the slope and y-intercept, we need to solve for y:
10x = 3y+6 OR y = (10/3)x - 2 slope m = 10/3, y-intercept (0,-2)
(0, -2) (3/5,0)